# Almost mathematics of pointed symmetric monoidal model categories by Smith ideal theory

### v4

@@ Abstract.
This article is a generalization of a result in Quillen’s note [Qui96] “Module theory over non-unital rings” giving a one-to-one correspondence between bilocalization of abelian categories of modules and idempotent ideals of the base ring. Faltings [Fal88]; Gabber and Ramero [GR03] established almost mathematics, the same as Quillen’s bilocalization of a category of modules by nil-modules.
In this paper, by using the theory of Smith ideals mentioned in Hovey [Hov14], we consider almost mathematics of symmetric monoidal pointed model categories and prove a weak analogue of the one-to-one correspondence in Quillen’s note [Qui96].
@@
Symmetric monoidal model categories, Smith ideals, almost mathematics, Bousfield localization and colocalization
@@ 1. Introduction
Faltings [Fal88] first introduced one of the topics of algebra theory almost mathematics. Later, Gabber and Ramero provided a textbook [GR03] of detailed foundation of almost mathematics, where they called those materials of the theory almost ring theory, almost modules, almost algebras, and almost homotopy algebras. Referring Gabber and Ramero’s book, almost mathematics is obtained by a localization of an abelian category of modules by almost isomorphisms: Let $V$ be a unital commutative ring with an idempotent ideal $\mathcal{m}$ of $V$. A $V$-module $M$ is said to be almost zero if $M$ is killed by $\mathcal{m}$, and almost mathematics is a theory of algebra working by localizing the category of $V$-modules by the Serre subcategory of almost zero modules. Almost isomorphisms are morphisms whose kernels and cokernels are both almost zero. In the localized category, module objects are called almost modules and algebra objects are called almost algebras.
Independently, Quillen considered the same theory of almost mathematics in unpublished note [Qui96] as linear algebra over non-unital rings, which he characterized as bilocalization - both localization and colocalization - of abelian categories of modules. Further, Quillen proved the following one-one correspondence between idempotent ideals and reflexive bilocalization functors.
@@
([Qui96] Proposition 6.5) Let $V$ be a commutative ring with a multiplicative unit. There is a one-to-one correspondence between Serre subcategories $\mathcal{S}$ of $\mathrm{Mod}_{V}$ that the localization $F:\mathrm{Mod}_{V}\to\mathrm{Mod}_{V}/\mathcal{S}$ is also a colocalization, and idempotent ideals of $V$. $\Box$
The purpose of this article is to establish an almost mathematics theory of symmetric monoidal model categories, where we prove a monoidal categorical analogue of Theorem 1.1. Namely, by utilizing the commutative Smith ideal theory developed by Hovey [Hov14] and White–Yau [WY24], we introduce homotopy analogues of idempotent ideals and almost isomorphisms, which we refer to as homotopically idempotent Smith ideals and almost weak equivalences. The main results of this paper are as follows, as mentioned in Theorem 2.11.
@@
Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, $j:I\to V$ be a homotopically idempotent commutative Smith ideal. Then the internal-hom functor
$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-):\mathcal{M}\to\mathcal{M}$
is equivalent to the Bousfield bilocalization functor by almost weak equivalences.
Since $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is a right Quillen functor, this theorem shows us that the Bousfield localization by almost weak equivalences is both a left and right Quillen functor.
Conversely, in the case that the monoidal unit generates the stable symmetric monoidal model category in the sense of MacLane [Mac88, Chapter V, Section 7, p.127], we can construct a homotopically idempotent ideal (Proposition 4.5 and Theorem 4.7).
Furthermore, we generalize some result of almost algebras for model category setting by using Smith ideal language. Again, let $V$ be a commutative ring with an idempotent ideal $\mathfrak{m}$. We assume that $\tilde{\mathfrak{m}}$ is $V$-flat. Let $\mathrm{CAlg}(\mathrm{Mod}_{V})$ denote the category of $V$-algebras and $\mathrm{CAlg}(\mathrm{Mod}_{V}^{\rm al})$ the category of almost algebras. Gabber and Ramero defined the functor
@@
for any almost unital $V$-algebra $A$ in [GR03, Section 2.2]. Those morphisms $\tilde{\mathfrak{m}}\to V\to A$ and $\tilde{\mathfrak{m}}\to\tilde{\mathfrak{m}}\otimes_{V}A\to A$ induce a ring homomorphism $A_{!!}\to A$, which is almost isomorphism. Due to [GR03, Proposition 2.2.29], the almost unitalization $(-)_{!!}$ is a left adjoint of the localization $\mathrm{CAlg}(\mathrm{Mod}_{V})\to\mathrm{CAlg}(\mathrm{Mod}_{V}^{\rm al})$. We call the left adjoint $(-)_{!!}$ almost unitalization in this paper. By using Smith ideal theory, we generalize the almost unitalization functor $(-)_{!!}$ that is right adjoint of the localization functor.
@@
In Section 3, we generalize the above almost unitalization functor by Smith ideal theory language, and we prove that it is a left adjoint of the Bousfield localization functor by the class of almost weak equivalences defined in Section 2, as Theorem 3.6 and Corollary 3.7.
In Section 4, as the inverse direction of Theorem 2.11, we demonstrate constructing homotopically idempotent Smith ideal from given symmetric monoidal bilocalization functor on symmetric monoidal model categories under some mild conditions, and the result is proven as Theorem 4.7.
Finally, in Section 5, we remark that the cokernel-and-kernel functors in the (operadic) Smith ideal theory induces a categorical equivalence between non-unital algebra objects and augmented algebra objects of stable symmetric monoidal model categories.
@@ 1.1. Example: The almost mathematics of a stable model category of chain complexes
In this section, we consider the model category of chain complexes of modules as an example of almost mathematics of model categories. We fix a commutative base ring $V$ and an idempotent ideal $\mathfrak{m}$ ($\mathfrak{m}^{2}=\mathfrak{m}$) of $V$, and we assume that the tensor product $\tilde{\mathfrak{m}}=\mathfrak{m}\otimes_{V}\mathfrak{m}$ is a flat $V$-module. Let $\operatorname{Ch}(V)$ denote the category of chain complexes of $V$-modules. Then $\operatorname{Ch}(V)$ has a canonical model structure called the injective model structure that cofibrations are degree-wise injections and weak equivalences are quasi-isomorphisms. Since we assume that $\tilde{\mathfrak{m}}$ is a flat $V$-module, the functor $\tilde{\mathfrak{m}}\otimes_{V}(-):\operatorname{Ch}(V)\to\operatorname{Ch}(V)$ is exact and a left Quillen functor, whose right adjoint is the mapping complex functor $\mathrm{Map}_{V}(\tilde{\mathfrak{m}},\,-)$.
@@
A perfectoid commutative valuation ring is defined to be a valuation ring of a perfectoid field defined in Scholze’s paper [Sch12, p.246, Definition 1.2]. Let $V$ be a perfectoid valuation ring. Then the set of topologically nilpotent elements $\mathfrak{m}$ of $V$ is an idempotent maximal ideal and a flat $V$-module, explained in, for example, Bhatt’s lecture note [Bha17, p.18, Example 4.1.3]. In general, the product map $\tilde{\mathfrak{m}}\to\mathfrak{m}$ is an isomorphism if $\mathfrak{m}$ is a flat $V$-module.
In the category $\mathrm{Ch}(V)$ of almost weak equivalences is defined as the following:
@@
A morphism $f:M\to N$ of $V$-complexes is said to be an almost quasi-isomorphism if the induced morphism $\tilde{\mathfrak{m}}\otimes_{V}f:\tilde{\mathfrak{m}}\otimes_{V}M\to\tilde{\mathfrak{m}}\otimes_{V}N$ is a quasi-isomorphism. Let $\operatorname{Ch}(V)^{\rm al}$ denote the Bousfield localization of $\operatorname{Ch}(V)$ by the class of almost quasi-isomorphisms. A chain complex $E$ is said to be almost local if $E$ is a local object with respect to almost quasi-isomorphisms. That is, for any almost quasi-isomorphism $f:M\to N$, the induced map
@@
In Quillen’s note [Qui96], a $V$-module $E$ that the induced homomorphism $E\to\mathrm{Hom}_{V}(\mathfrak{m},\,E)$ is an isomorphism is said to be closed, and closed modules are almost local objects and vice versa. The following lemma characterizes almost local objects.
@@
Conversely, assume that $E$ is almost local. Since the canonical morphism $\tilde{\mathfrak{m}}\otimes_{V}M\to M$ is an almost quasi-isomorphism for any $M\in\operatorname{Ch}(V)$, then the induced morphism
@@
Then the mapping complex functor
@@
Acknowledgments
I would like to thank anonymous referees for giving me innovative ideas and constructive comments to use Hovey’s Smith ideal theory and White and Yau’s operadic Smith ideal theory.
@@ 2. Almost mathematics of pointed symmetric monoidal model categories
In this section, referring White and Yau’s work [WY24], we recall definition and some properties of symmetric monoidal model model categories and (commutative) Smith ideals of them. Next, in order to establish almost mathematics, we define homotopically idempotent Smith ideals, and we prove this paper’s main result Theorem 2.11.
@@
${\mathcal{M}^{\rm cof}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{M}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R}$$\scriptstyle{Q}$${\mathcal{M}^{\rm fib}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
@@ 2.1. Commutative Smith ideals of symmetric monoidal model categories
A symmetric monoidal model category $\mathcal{M}$ is a model category with a symmetric monoidal structure $-\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ satisfying the following two axioms:
@@
We say that the symmetric monoidal structure on $\mathcal{M}$ is closed if $\mathcal{M}$ contains an internal hom-object $\mathrm{Map}_{\mathcal{M}}(M,\,N)$ with respect to the monoidal structure $-\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ for any $M,\,N\in\mathcal{M}$. In this case, the push-out product axiom implies that, for any cofibrant object $M$, both those induced functors $-\otimes M:\mathcal{M}\to\mathcal{M}$ and $M\otimes-:\mathcal{M}\to\mathcal{M}$ are left Quillen functors.
Further, we assume that all model categories in this paper are cofibrantly generated. For any class $\mathcal{T}$ of morphisms in a category, ${}^{\boxslash}\mathcal{T}$ (resp. $\mathcal{T}^{\boxslash}$) denotes the class of morphisms having the right (resp. left) lifting property with respect to all morphisms in $\mathcal{T}$. Then ${}^{\boxslash}(\mathcal{T}^{\boxslash})$ is said to be the weakly saturated class of morphisms generated by $\mathcal{T}$. In this paper, we always assume that all model categories are cofibrantly generated: There exist small subsets $\mathcal{I}$ and $\mathcal{J}$ of the set of morphisms of the model category such that the collection of all cofibrations (resp. trivial cofibrations) is the weakly saturated class of morphisms generated by $\mathcal{I}$ (resp. $\mathcal{J}$). In this paper, we call $\mathcal{I}$ (resp. $\mathcal{J}$) the set of generating cofibrations (resp. the set of generating trivial cofibrations). Further, if a cofibrantly generated model category is locally presentable, it is said to be combinatorial. By the theorem of Barwick and Smith [Bar10, Theorem 4.7], for any left proper and combinatorial model category and a small set of homotopy class of morphisms of it, the Bousfield localization exists such that the localized model category is also left proper and combinatorial.
Let $\Delta^{1}$ denote the category with two objects 00 and $1$, and only one non-trivial morphism $0\to 1$. The functor category $\mathrm{Fun}(\Delta^{1},\,\mathcal{M})$ from $\Delta^{1}$ to $\mathcal{M}$ is called the arrow category of $\mathcal{M}$, and $\mathrm{Ar}(\mathcal{M})$ denotes it.
By Hovey [Hov14, Theorem 1.2], the arrow category $\mathrm{Ar}(\mathcal{M})$ has two distinct symmetric monoidal structures derived from $\mathcal{M}$’s: For any two morphisms in $\mathcal{M}$, $f:X_{0}\to X_{1}$ and $g:Y_{0}\to Y_{1}$, one has a commutative square:
@@
One of them is the diagonal (or tensor) product monoidal structure is defined by $f\otimes g$ as the composition
@@
[…]
@@
Those monoidal model categories $\mathrm{Ar}^{\Box}(\mathcal{M})$ and $\mathrm{Ar}^{\otimes}(\mathcal{M})$ are connected the cokernel-and-kernel adjunction
@@
which is a Quillen adjunction. The cokernel functor $\mathrm{cok}$ is defined by the sending
@@
respectively. In the case that $\mathcal{M}$ is a pointed, left proper, and combinatorial symmetric monoidal model category, the adjunction is lax monoidal Quillen by Hovey’s work (the proof is available at [WY24, Proposition 2.4.3]). Further, if $\mathcal{M}$ is stable, in addition, the adjunction $(\mathrm{cok},\,\mathrm{ker})$ is a Quillen equivalence. (See Hovey [Hov14, Theorem 1.4] or White and Yau [WY24, Proposition 2.3]).
We recall the definition of commutative Smith ideals. A commutative monoid object $A$ of $\mathcal{M}$ is a monoid whose product $\mu_{A}:A\otimes A\to A$ commutes the braiding $\tau_{A,\,A}:A\otimes A\to A\otimes A$, denoting $\mathrm{CAlg}(\mathcal{M})$ the subcategory of commutative objects of $\mathcal{M}$.
Definition 2.1 (c.f. [WY24] Definition 3.3.1).

### v5

@@ Abstract.
This article is a generalization of a result in Quillen’s note [14] “Module theory over non-unital rings” giving a one-to-one correspondence between bilocalization of abelian categories of modules and idempotent ideals of the base ring. Faltings [4]; Gabber and Ramero [5] established almost mathematics, the same as Quillen’s bilocalization of a category of modules by nil-modules.
In this paper, by using the theory of Smith ideals mentioned in Hovey [9], we consider almost mathematics of symmetric monoidal pointed model categories and prove a weak analogue of the one-to-one correspondence in Quillen’s note [14].
@@
Almost mathematics, Smith ideals, monoidal model categories, Bousfield localization and colocalization
1991 Mathematics Subject Classification:
18N55 (primary), 18N70 (secondary)
The author was supported by Grants-in-Aid for Scientific Research No. 23K030080, Japan Society for the Promotion of Science.
@@ 1. Introduction
Faltings [4] first introduced the theory of almost mathematics to apply $p$-adic Hodge theory, and subsequently, Gabber and Ramero [5] developed the theory of algebra within this framework. In their terminology, the theory includes almost ring theory, almost modules, almost algebras, and almost homotopy algebras. Building on Gabber and Ramero’s work, almost mathematics is formulated by localizing an abelian category of modules at almost isomorphisms: let $V$ be a unital commutative ring with an idempotent ideal $\mathfrak{m}$ of $V$. A $V$-module $M$ is defined to be almost zero if it is annihilated by $\mathfrak{m}$, and almost mathematics is an algebraic theory constructed by localizing the category of $V$-modules with respect to the Serre subcategory of almost zero modules. Almost isomorphisms are morphisms whose kernels and cokernels are both almost zero. In the localized category, module objects are referred to as almost modules , while algebra objects are called almost algebras.
Independently, in an unpublished note [14], Quillen considered the same theory of almost mathematics as linear algebra over non-unital rings. He characterized it as a bilocalization, that is, both a localization and a colocalization, of abelian categories of modules. Further, Quillen proved the following one-to-one correspondence between idempotent ideals and reflexive bilocalization functors.
@@
([14] Proposition 6.5) Let $V$ be a commutative ring with a multiplicative unit. There is a one-to-one correspondence between idempotent ideals of $V$ and Serre subcategories $\mathcal{S}$ of $\mathrm{Mod}_{V}$ for which the localization $F:\mathrm{Mod}_{V}\to\mathrm{Mod}_{V}/\mathcal{S}$ is also a colocalization. $\Box$
The purpose of this article is to establish a theory of almost mathematics for symmetric monoidal model categories and to prove a monoidal categorical analogue of Theorem 1.1. Namely, using the commutative Smith ideal theory developed by Hovey [9] and White–Yau [18], we introduce homotopy analogues of idempotent ideals and almost isomorphisms, which we refer to as homotopically idempotent Smith ideals and almost weak equivalences. The main result of this paper are summarized in Theorem 2.12.
@@
Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, and let $j:I\to V$ be a homotopically idempotent commutative Smith ideal. Then the internal-hom functor
[…]
is equivalent to the Bousfield bilocalization functor with respect to almost weak equivalences.
[…]
Conversely, in the case that the monoidal unit generates the stable symmetric monoidal model category in the sense of MacLane [11, Chapter V, Section 7, p.127], we can construct a homotopically idempotent ideal (Proposition 4.5 and Theorem 4.7).
Furthermore, we generalize some results on almost algebras to the model category setting using the language of Smith ideals. Again, let $V$ be a commutative ring with an idempotent ideal $\mathfrak{m}$. We assume that $\tilde{\mathfrak{m}}$ is $V$-flat. Let $\mathrm{CAlg}(\mathrm{Mod}_{V})$ denote the category of $V$-algebras and $\mathrm{CAlg}(\mathrm{Mod}_{V}^{\rm al})$ the category of almost algebras. Gabber and Ramero defined the functor
@@
for any almost unital $V$-algebra AA in [5, Section 2.2]. Those morphisms $\tilde{\mathfrak{m}}\to V\to A$ and $\tilde{\mathfrak{m}}\to\tilde{\mathfrak{m}}\otimes_{V}A\to A$ induce a ring homomorphism $A_{!!}\to A$, which is an almost isomorphism. Due to [5, Proposition 2.2.29], the almost unitalization $(-)_{!!}$ is a left adjoint of the localization $\mathrm{CAlg}(\mathrm{Mod}_{V})\to\mathrm{CAlg}(\mathrm{Mod}_{V}^{\rm al})$. We call the left adjoint $(-)_{!!}$ almost unitalization in this paper. Using Smith ideal theory, we generalize the almost unitalization functor $(-)_{!!}$ in the language of Smith ideals.
@@
In Section 3, we generalize the above almost unitalization functor in the language of Smith ideal theory, and we prove that it is a left adjoint of the Bousfield localization functor with respect to the class of almost weak equivalences defined in Section 2; see Theorem 3.6 and Corollary 3.7.
In Section 4, as the inverse direction of Theorem 2.12, we construct a homotopically idempotent Smith ideal from a given symmetric monoidal bilocalization functor on a symmetric monoidal model category under mild conditions. The result is proved as Theorem 4.7.
Finally, in Section 5, we remark that the cokernel-and-kernel functors in the (operadic) Smith ideal theory induce a categorical equivalence between non-unital algebra objects and augmented algebra objects of stable symmetric monoidal model categories.
For the reader’s convenience, we mention related independent work of Hebestreit and Scholze on higher almost ring theory [7]. Their work develops a higher almost mathematics framework in an $\infty$-categorical setting. The present paper is complementary, developing a model-categorical framework based on Smith ideal techniques. To clarify the independent development of the two projects, we also record that the fourth version of the present preprint appeared as arXiv:2301.04541v4 on July 29, 2024 UTC (July 30 in Japan), and that the first version of Hebestreit and Scholze’s preprint appeared as arXiv:2409.01940v1 on September 3, 2024.
@@ 1.1. Example: The almost mathematics of a stable model category of chain complexes
[…]
@@
A perfectoid commutative valuation ring is a valuation ring of a perfectoid field in the sense of Scholze [15, p.246, Definition 1.2]. Let $V$ be a perfectoid valuation ring. Then the set of topologically nilpotent elements $\mathfrak{m}$ of $V$ is an idempotent maximal ideal and a flat $V$-module, as explained, for example, in Bhatt’s lecture note [3, p.18, Example 4.1.3]. In general, the product map $\tilde{\mathfrak{m}}\to\mathfrak{m}$ is an isomorphism if $\mathfrak{m}$ is a flat $V$-module.
In the category $\mathrm{Ch}(V)$, almost weak equivalences are defined as follows:
@@
A morphism $f:M\to N$ of $V$-complexes is said to be an almost quasi-isomorphism if the induced morphism $\tilde{\mathfrak{m}}\otimes_{V}f:\tilde{\mathfrak{m}}\otimes_{V}M\to\tilde{\mathfrak{m}}\otimes_{V}N$ is a quasi-isomorphism. Let $\operatorname{Ch}(V)^{\rm al}$ denote the Bousfield localization of the model category $\operatorname{Ch}(V)$ by the class of almost quasi-isomorphisms. A chain complex $E$ is said to be almost local if $E$ is a local object with respect to almost quasi-isomorphisms. That is, for any almost quasi-isomorphism $f:M\to N$, the induced map
@@
In Quillen’s note [14], a $V$-module $E$ is said to be closed if the induced homomorphism $E\to\mathrm{Hom}_{V}(\mathfrak{m},\,E)$ is an isomorphism, and it is shown that closed modules are almost local objects and vice versa; see [14, Proposition 5.3]. The following lemma characterizes almost local objects.
@@
Conversely, assume that $E$ is almost local. Since the canonical morphism $\tilde{\mathfrak{m}}\otimes_{V}M\to M$ is an almost quasi-isomorphism for any $M\in\operatorname{Ch}(V)$, the induced morphism
@@
The mapping complex functor
@@ 2. Almost mathematics of pointed symmetric monoidal model categories
In this section, following White and Yau’s work [18], we recall the definition and some properties of symmetric monoidal model categories and their (commutative) Smith ideals. Next, in order to establish almost mathematics, we define homotopically idempotent Smith ideals, and we prove the main result, Theorem 2.12.
A category is pointed if it has an initial object $\emptyset$ and a final object * such that the unique morphism $\emptyset\to*$ is an isomorphism. We call the initial object the zero object, and 0 denotes it. Throughout this paper, we consider pointed (model) categories.
@@ 2.1. Commutative Smith ideals of symmetric monoidal model categories
A symmetric monoidal model category $\mathcal{M}$ is a model category with a symmetric monoidal structure $\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ satisfying the following two axioms:
@@
We say that the symmetric monoidal structure on $\mathcal{M}$ is closed if $\mathcal{M}$ contains an internal hom-object $\mathrm{Map}_{\mathcal{M}}(M,\,N)$ with respect to the monoidal structure $\otimes-:\mathcal{M}\times\mathcal{M}\to\mathcal{M}$ for any $M,\,N\in\mathcal{M}$. In this case, the push-out product axiom implies that, for any cofibrant object $M$, both induced functors $\otimes M:\mathcal{M}\to\mathcal{M}$ and $M\otimes-:\mathcal{M}\to\mathcal{M}$ are left Quillen functors.
Further, we assume that all model categories in this paper are cofibrantly generated. For any class $\mathcal{T}$ of morphisms in a category, $^{\boxslash}\mathcal{T}$ (resp. $\mathcal{T}^{\boxslash}$) denotes the class of morphisms having the right (resp. left) lifting property with respect to all morphisms in $\mathcal{T}$. Then $^{\boxslash}(\mathcal{T}^{\boxslash})$ is said to be the weakly saturated class of morphisms generated by $\mathcal{T}$. In this paper, we always assume that all model categories are cofibrantly generated: There exist small subsets $\mathcal{I}$ and $\mathcal{J}$ of the set of morphisms of the model category such that the collection of all cofibrations (resp. trivial cofibrations) is the weakly saturated class of morphisms generated by $\mathcal{I}$ (resp. $\mathcal{J}$). In this paper, we call $\mathcal{I}$ (resp. $\mathcal{J}$) the set of generating cofibrations (resp. the set of generating trivial cofibrations). Further, if a cofibrantly generated model category is locally presentable, it is said to be combinatorial. By the theorem of Barwick and Smith [2, Theorem 4.7], for any left proper and combinatorial model category and a small set of homotopy classes of morphisms in it, the Bousfield localization exists such that the localized model category is also left proper and combinatorial.
Let $\Delta^{1}$ denote the category with two objects 0 and $1$, and only one non-trivial morphism $0\to 1$. The functor category $\mathrm{Fun}(\Delta^{1},\,\mathcal{M})$ from $\Delta^{1}$ to $\mathcal{M}$ is called the arrow category of $\mathcal{M}$ and is denoted by $\mathrm{Ar}(\mathcal{M})$.
By Hovey [9, Theorem 1.2], the arrow category $\mathrm{Ar}(\mathcal{M})$ has two distinct symmetric monoidal structures derived from that of $\mathcal{M}$. For any two morphisms $f:X_{0}\to X_{1}$ and $g:Y_{0}\to Y_{1}$ in $\mathcal{M}$, one has a commutative square:
@@
The diagonal (or tensor) product monoidal structure is defined by $f\otimes g$ as the composition
@@
The arrow category $\mathrm{Ar}(\mathcal{M})$ has the evaluation functors $\mathrm{Ev}_{i}(f:X_{0}\to X_{1})=X_{i}\ (i=0,\,1)$, which have a left adjoint and a right adjoint:
@@
These monoidal model categories $\mathrm{Ar}^{\Box}(\mathcal{M})$ and $\mathrm{Ar}^{\otimes}(\mathcal{M})$ are connected by the cokernel-and-kernel adjunction
@@
which is a Quillen adjunction. The cokernel functor $\mathrm{cok}$ is defined by
@@
respectively. In the case that $\mathcal{M}$ is a pointed, left proper, and combinatorial symmetric monoidal model category, the adjunction is lax monoidal Quillen by Hovey’s work (the proof is available at [18, Proposition 2.4.3]). Further, if $\mathcal{M}$ is stable, in addition, the adjunction $(\mathrm{cok},\,\mathrm{ker})$ is a Quillen equivalence. (See Hovey [9, Theorem 1.4] or White and Yau [18, Proposition 2.3]).
We recall the definition of commutative Smith ideals. A commutative monoid object $A$ of $\mathcal{M}$ is a monoid whose product $\mu_{A}:A\otimes A\to A$ commutes with the braiding $\tau_{A,\,A}:A\otimes A\to A\otimes A$. We denote by $\mathrm{CAlg}(\mathcal{M})$ the subcategory of commutative objects of $\mathcal{M}$.
Definition 2.1 (c.f. [18] Definition 3.3.1).